# Massive-STEPS: Massive Semantic Trajectories for Understanding POI Check-ins – Dataset and Benchmarks

### v1

@@ 1 Introduction
Human mobility data and modeling are essential for understanding how individuals interact with and move through physical spaces. This understanding enables a wide range of applications, including urban planning yuan2025worldmove , travel service recommendations feng-etal-2025-agentmove , improved commercial advertising strategies 10.1145/3477495.3531983 , and Point-of-Interest (POI) recommendation 10.1145/3369822 ; 10.1145/3626772.3657840 ; zhang2025survey . Recently, human mobility data has become even more crucial with the increasing use of large language models (LLM) agents to simulate human-like behavior and routines zhou2024large ; jiawei2024large . However, while simulated human mobility data are starting to gain popularity 10.1145/3624475 ; 10.1145/3394486.3412862 , they may not accurately reflect real-world human behavior SALIM2020106964 , highlighting the value of evaluating on real-world data. These advancements are enabled by and large with the rise of Location-based Social Networks (LBSNs), which generate vast amounts of spatio-temporal data through user check-ins  zhang2025survey ; 10.1145/3626772.3657840 . This rich data source has allowed the development of POI recommendation systems that leverage users’ historical visiting behaviors to suggest relevant locations. Such systems enhance user engagement through personalization and provide commercial value to both users and businesses by aligning recommendations with individual preferences and available services zhang2025survey ; 10.1145/3369822 .
@@
Our paper addresses three critical gaps in POI recommendation research and datasets, as detailed in Section 2.1. First, as shown in Fig. 1, the field remains dominated by studies focused on just two cities, New York and Tokyo, based on the Foursquare dataset curated in yang2014modeling . This dataset, collected in 2012-2013, raises concerns about its temporal quality, as many POIs may no longer exist and user behavior may have changed yeow2021point . While some recent studies have expanded to other cities 10.1007/s11280-024-01279-y ; merinov2024positive ; feng-etal-2025-agentmove , they often rely on the Global-scale Check-in Dataset (GSCD) yang2015nationtelescope ; yang2016participatory , which, despite its large coverage, is also from 2012-2013 and contains nearly 50% erroneous entries monti2018semantic .
Second, most existing studies are difficult to reproduce, either due to the lack of clearly defined geographic boundaries or the unavailability of the datasets themselves, hindering fair comparison and replication. Finally, we join recent efforts yuan2025worldmove in advocating for the inclusion of low-resource and underrepresented cities in evaluation. Expanding beyond well-studied urban centers is essential for building more generalizable and universally applicable POI recommendation models. Table 1 summarizes these limitations in terms of geographic coverage, temporal span, and reproducibility.
Table 1: Comparison of check-in datasets commonly used for POI recommendation tasks. GSCD  yang2014modeling ; yang2016participatory and Semantic Trails monti2018semantic are global datasets not grouped into individual cities, whereas others perform city-level grouping. Replicable indicates whether city boundaries are clearly defined or can be reliably reconstructed.
@@
GSCD yang2014modeling ; yang2016participatory Varies 2012-2013 17    Coordinates, Category N/A
Semantic Trails monti2018semantic Varies 2012-2013, 2017-2018 24    Category N/A
NYC and Tokyo yang2014modeling 2 2012-2013 11    Coordinates, Category
Gowalla-CA cho2011friendship ; 10.1145/2484028.2484030 1 2009-2010 21    Coordinates, Category
AgentMove feng-etal-2025-agentmove 12 2012-2013 17    Coordinates, Category
@@
In this paper, we introduced the Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) Dataset, derived from the Semantic Trails dataset (STD) monti2018semantic . Massive-STEPS includes high-quality check-ins from 2012-2013 and 2017-2018, providing more modern and updated POI check-in data. This supports longitudinal POI recommendation studies and addresses the limitations of older datasets commonly used in prior studies. The dataset covers 12 diverse cities across multiple regions, including East, West, and Southeast Asia, North and South America, Australia, the Middle East, and Europe. Notably, we placed a deliberate emphasis on under-explored regions by including cities such as Jakarta, Kuwait City, and Petaling Jaya, filling a key gap in POI recommendation research that has largely focused on major urban centers. We further enriched STD by aligning it with Foursquare’s Open Source Places dataset, incorporating metadata such as POI coordinates, POI names, and addresses—details unavailable in the original STD.
@@ 2 Related Works
Figure 1: Distribution of POI recommendation studies modeled on specific cities, filtered from Table IV in zhang2025survey . We identified and counted studies that explicitly mentioned city names, revealing the skewness of existing research, which is saturated around New York and Tokyo. In addition, we include the distribution of studies by LBSN platform, showing that Foursquare is by far the most commonly used source of check-in data. The list of identified studies is shown in Table 5.
@@ 2.1 Existing Datasets
A survey conducted by zhang2025survey outlines the broad landscape of POI recommendation research, covering a wide range of models and architectures used in prior studies. While it offers a high-level overview of the datasets used, it lacks a dedicated discussion or evaluation of POI datasets. We address this gap by analyzing commonly used datasets in the field and positioning our dataset within this context.
@@
Building on the tabular summary provided in zhang2025survey , which offers a representative overview of the broader POI recommendation literature, we investigated which datasets are most commonly used in prior studies. From their original table (Table IV), we filtered entries pertaining specifically to POI and next POI recommendation tasks and identified (1) the most frequently used LBSN check-in data sources and (2) the most commonly studied cities. As shown in Fig. 1, Foursquare remains the dominant source of LBSN data in existing studies, appearing in almost 50% of the surveyed works. While several variants of Foursquare datasets have been employed, the most widely used are the NYC and Tokyo Dataset yang2014modeling (often abbreviated as FSQ-NYC and FSQ-TKY) and the Global-scale Check-in Dataset (GSCD) yang2015nationtelescope ; yang2016participatory , curated by the same authors. Other LBSN sources occasionally used include Gowalla cho2011friendship , Brightkite cho2011friendship , and Weeplaces 10.14778/3115404.3115407 .
@@
Due to the widespread use of FSQ-NYC and FSQ-TKY yang2014modeling , the majority of POI recommendation studies are disproportionately focused on these two cities, as illustrated in Fig. 1. While there is nothing inherently problematic about studying NYC and Tokyo, there has been growing interest in expanding research to a broader range of cities, particularly those that are underexplored or considered low-resource yuan2025worldmove , as cultural and regional differences influence collective mobility behaviors. For instance, in some urban contexts, residents tend to commute to central business districts in the morning, whereas in others, nightlife activities such as visiting bars after work are more common yang2015nationtelescope . Ensuring diverse geographic coverage is increasingly important, especially as LLMs are adopted for POI recommendation tasks. LLMs are known to exhibit geographical biases, often underperforming in regions with lower socioeconomic conditions manvi2024large . It remains an open question whether LLM-based POI recommendation models can effectively generalize across diverse urban environments.
In addition, because many studies rely on the FSQ-NYC and FSQ-TKY, they are often constrained to the timespan it covers: check-in data from 2012 to 2013. However, POI data is inherently dynamic: venues may have closed, relocated, or changed in category over time. yeow2021point underscores the importance of validating the temporal quality of POI datasets by recording whether and when a venue’s information has been updated to reflect real-world changes. This is particularly critical, as recommender systems should avoid suggesting POIs that no longer exist or have undergone substantial changes (e.g., a former bookstore converted into a coworking space) and behave dynamically over longitudinal periods Yabe2024 . Moreover, behavioral patterns captured over a decade ago may no longer align with modern user preferences and routines. For example, the opening of a new train station may significantly shift commuting patterns and the popularity of surrounding POIs.
@@
More recently, researchers have begun leveraging the broader Global-scale Check-in Dataset (GSCD) yang2015nationtelescope ; yang2016participatory , which spans 415 cities across 77 countries. Despite its wider geographic coverage, GSCD is temporally limited to the same 2012-2013 period as FSQ-NYC and FSQ-TKY, and thus suffers from similar issues of temporal quality. More critically, monti2018semantic demonstrated that GSCD suffers from significant data quality issues, with over 14 million check-ins (about 44%) of the dataset flagged as erroneous due to anomalous user behavior. These include (1) repeated check-ins at the same venue, (2) check-ins occurring within implausibly short time intervals (less than one minute), and (3) transitions between venues that would require travel speeds exceeding Mach 1, which are physically unreasonable.
To address these limitations, monti2018semantic introduced the Semantic Trails Dataset (STD), which applies systematic filtering procedures to enhance data quality. STD comprises two subsets: a cleaned version of GSCD covering 2012-2013 (STD 2013), and a newer collection of check-ins from 2017-2018 (STD 2018), sourced from Foursquare Swarm. STD 2018 also spans a wider range of cities, making it valuable for capturing globally distributed user behavior, in contrast to GSCD’s focus on densely populated urban centers. Both subsets follow the same rigorous filtering criteria, resulting in a higher-quality check-in dataset for POI recommendation tasks. Given these improvements, we adopted STD as the source for our check-in dataset.
@@
Another persistent challenge in POI recommendation research is the lack of reproducibility in dataset preprocessing. While some recent studies utilize datasets like GSCD to cover a wide range of cities, they often omit important details needed for replicating their data filtering processes. For example, AgentMove feng-etal-2025-agentmove conducted city-level filtering based on a minimum distance between each trajectory and the city center, but they did not specify how the city center was defined or what distance threshold was used. Generally, city filtering can be done in two ways: (1) by selecting points within a fixed radius from a central coordinate, or (2) by applying precise administrative boundaries, such as those from OpenStreetMap. The radius-based approach is less reliable, as it assumes a circular city shape and requires a subjective definition of the city’s center. In contrast, the boundary-based method is more robust and reproducible, and is the approach we adopted in this work. Without standardized, transparent filtering criteria and publicly available datasets, it remains difficult to ensure fair comparisons across POI recommendation methods.
@@ 2.2 Understanding Urban Features and POI Recommendation
POI recommendation studies that evaluate models across multiple city-level datasets often include analyses to assess how well their methods generalize across different urban contexts. It is well understood that POI recommendation accuracy metrics (e.g., Acc@k, NDCG@k) can vary substantially between cities and can be interpreted as a proxy for how easy or difficult a city is to model. The assumption is that higher performance reflects more predictable or structured mobility patterns. This viewpoint is consistent with prior work highlighting the role of cultural and urban-specific factors in shaping mobility behaviors yang2015nationtelescope ; sun2024city .
Several studies have proposed hypotheses connecting specific urban features to modeling difficulty. For example, GETNext yang2022getnext hypothesized that cities with fewer check-ins and higher spatial sparsity of POIs are harder to model. STHGCN 10.1145/3539618.3591770 suggested that a larger number of user trajectories improves predictive accuracy by providing richer collaborative signals, whose architecture is designed to leverage. LLM4POI 10.1145/3626772.3657840 proposed that cities with a greater variety of POI categories are easier to model due to LLMs’ contextual reasoning capabilities, whereas cities covering a broader geographic area tend to be more difficult to model. In the zero-shot POI recommendation setting, AgentMove  feng-etal-2025-agentmove reported two key findings: (1) geospatial biases inherent in LLMs can hinder prediction quality across cities, and (2) LLMs are influenced by city-specific mobility patterns.
@@ 3 Dataset
Table 2: Summary statistics of the 12 Massive-STEPS subsets, including the number of users, trajectories, POI locations, total check-ins, and train, validation, and test sample counts. For comparison, we also include statistics from existing Foursquare-based yang2014modeling and Gowalla-based cho2011friendship datasets. Due to variations in dataset preprocessing across studies, we report the version used in 10.1145/3539618.3591770 .
@@
NYC and Tokyo Check-in Dataset yang2014modeling
@@ 3.1 Creation Process
Massive-STEPS is derived from the two subsets of STD monti2018semantic , incorporating check-ins from both the 2013 and 2018 subsets. We utilize two additional components from STD: (1) the cities metadata file, which provides the latitude and longitude of administrative regions (e.g., towns, suburbs) along with their corresponding country codes obtained from GeoNames; and (2) the POI category mapping, which links each Foursquare Category ID to its descriptive name (e.g., "Restaurant"). Based on this metadata, each POI is thus associated with several attributes: Foursquare Place ID, Foursquare Category ID, category name, latitude/longitude of the administrative region, the administrative region name, and the country code. For anonymization purposes and model training compatibility, we apply ordinal encoding to the Place IDs and Category IDs, assigning each a unique integer index.
@@ 3.1.1 Preprocessing
We split trajectories into training, validation, and test sets in a ratio of 7:1:2, following feng-etal-2025-agentmove . We ensure that all users in the test set appear at least once in the training or validation set, a common practice in prior studies yang2022getnext ; 10.1145/3539618.3591770 .
@@ 3.2 Description and Addressing Literature Gaps
Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) dataset is a city-level POI check-in dataset derived from Semantic Trails monti2018semantic , comprising user check-in trajectories from 12 cities: Beijing, Istanbul, Jakarta, Kuwait City, Melbourne, Moscow, New York, Petaling Jaya, São Paulo, Shanghai, Sydney, and Tokyo. It features anonymized POI check-ins enriched with geographical metadata to support spatiotemporal and sequential modeling tasks. City-level statistics, along with comparisons to existing datasets, are presented in Table 2. Fig. 2 shows a world map highlighting the locations of all cities included in the Massive-STEPS dataset. Table 6 shows the available fields in the dataset and provides an example for each field.
Massive-STEPS offers a more comprehensive and diverse representation of urban mobility compared to typical POI check-in datasets. As shown in Table 2, datasets like FSQ-NYC and FSQ-TKY yang2014modeling contain fewer than 10,000 candidate POI locations. In contrast, cities in Massive-STEPS cover significantly more POIs: Massive-STEPS New York has over 49,000 POIs, while Massive-STEPS Jakarta exceeds 76,000. Massive-STEPS Istanbul, one of the largest subsets, features a large user base of 23,700, offering a broad range of user behaviors. Although some Massive-STEPS subsets are smaller than their FSQ counterparts (e.g., Tokyo), we attribute this to the strict filtering procedures applied by STD to remove erroneous entries, as explained in Section 2.1. This scale introduces additional computational challenges. Models that rely on dense POI-to-POI adjacency matrices, for instance, may require more efficient implementations to avoid excessive memory consumption.

### v2

@@ 1 Introduction
Human mobility data and modeling are essential for understanding how individuals interact with and move through physical spaces. This understanding enables a wide range of applications, including urban planning [63], travel service recommendations [7], improved commercial advertising strategies [59], and Point-of-Interest (POI) recommendation [6, 22, 68]. Recently, human mobility data has become even more crucial with the increasing use of large language models (LLM) agents to simulate human-like behavior and routines [71, 20]. However, while simulated human mobility data are starting to gain popularity [31, 9], they may not accurately reflect real-world human behavior [33], highlighting the value of evaluating on real-world data. These advancements are enabled by and large with the rise of Location-based Social Networks (LBSNs), which generate vast amounts of spatio-temporal data through user check-ins  [68, 22]. This rich data source has allowed the development of POI recommendation systems that leverage users’ historical visiting behaviors to suggest relevant locations. Such systems enhance user engagement through personalization and provide commercial value to both users and businesses by aligning recommendations with individual preferences and available services [68, 6].
@@
Our paper addresses three critical gaps in POI recommendation research and datasets, as detailed in Section 2.1. First, as shown in Fig. 1, the field remains dominated by studies focused on just two cities, New York and Tokyo, based on the Foursquare dataset curated in [57]. This dataset, collected in 2012-2013, raises concerns about its temporal quality, as many POIs may no longer exist and user behavior may have changed [61]. While some recent studies have expanded to other cities [65, 28, 7], they often rely on the Global-scale Check-in Dataset (GSCD) [55, 56], which, despite its large coverage, is also from 2012-2013 and contains nearly 50% erroneous entries [29].
Second, most existing studies are difficult to reproduce, either due to the lack of clearly defined geographic boundaries or the unavailability of the datasets themselves, hindering fair comparison and replication. Finally, we join recent efforts [63] in advocating for the inclusion of low-resource and underrepresented cities in evaluation. Expanding beyond well-studied urban centers is essential for building more generalizable and universally applicable POI recommendation models. Table 1 summarizes these limitations in terms of geographic coverage, temporal span, and reproducibility.
Table 1: Comparison of check-in datasets commonly used for POI recommendation tasks. GSCD  [57, 56] and Semantic Trails [29] are global datasets not grouped into individual cities, whereas others perform city-level grouping. Replicable indicates whether city boundaries are clearly defined or can be reliably reconstructed.
@@
GSCD [57, 56] Varies 2012-2013 17    Coordinates, Category N/A
Semantic Trails [29] Varies 2012-2013, 2017-2018 24    Category N/A
NYC and Tokyo [57] 2 2012-2013 11    Coordinates, Category
Gowalla-CA [4, 62] 1 2009-2010 21    Coordinates, Category
AgentMove [7] 12 2012-2013 17    Coordinates, Category
@@
In this paper, we introduced the Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) Dataset, derived from the Semantic Trails dataset (STD) [29]. Massive-STEPS includes high-quality check-ins from 2012-2013 and 2017-2018, providing more modern and updated POI check-in data. This supports longitudinal POI recommendation studies and addresses the limitations of older datasets commonly used in prior studies. The dataset covers 12 diverse cities across multiple regions, including East, West, and Southeast Asia, North and South America, Australia, the Middle East, and Europe. Notably, we placed a deliberate emphasis on under-explored regions by including cities such as Jakarta, Kuwait City, and Petaling Jaya, filling a key gap in POI recommendation research that has largely focused on major urban centers. We further enriched STD by aligning it with Foursquare’s Open Source Places dataset, incorporating metadata such as POI coordinates, POI names, and addresses—details unavailable in the original STD.
@@ 2 Related Works
Figure 1: Distribution of POI recommendation studies modeled on specific cities, filtered from Table IV in [68]. We identified and counted studies that explicitly mentioned city names, revealing the skewness of existing research, which is saturated around New York and Tokyo. In addition, we include the distribution of studies by LBSN platform, showing that Foursquare is by far the most commonly used source of check-in data. The list of identified studies is shown in Table 5.
@@ 2.1 Existing Datasets
A survey conducted by [68] outlines the broad landscape of POI recommendation research, covering a wide range of models and architectures used in prior studies. While it offers a high-level overview of the datasets used, it lacks a dedicated discussion or evaluation of POI datasets. We address this gap by analyzing commonly used datasets in the field and positioning our dataset within this context.
@@
Building on the tabular summary provided in [68], which offers a representative overview of the broader POI recommendation literature, we investigated which datasets are most commonly used in prior studies. From their original table (Table IV), we filtered entries pertaining specifically to POI and next POI recommendation tasks and identified (1) the most frequently used LBSN check-in data sources and (2) the most commonly studied cities. As shown in Fig. 1, Foursquare remains the dominant source of LBSN data in existing studies, appearing in almost 50% of the surveyed works. While several variants of Foursquare datasets have been employed, the most widely used are the NYC and Tokyo Dataset [57] (often abbreviated as FSQ-NYC and FSQ-TKY) and the Global-scale Check-in Dataset (GSCD) [55, 56], curated by the same authors. Other LBSN sources occasionally used include Gowalla [4], Brightkite [4], and Weeplaces [26].
@@
Due to the widespread use of FSQ-NYC and FSQ-TKY [57], the majority of POI recommendation studies are disproportionately focused on these two cities, as illustrated in Fig. 1. While there is nothing inherently problematic about studying NYC and Tokyo, there has been growing interest in expanding research to a broader range of cities, particularly those that are underexplored or considered low-resource [63], as cultural and regional differences influence collective mobility behaviors. For instance, in some urban contexts, residents tend to commute to central business districts in the morning, whereas in others, nightlife activities such as visiting bars after work are more common [55]. Ensuring diverse geographic coverage is increasingly important, especially as LLMs are adopted for POI recommendation tasks. LLMs are known to exhibit geographical biases, often underperforming in regions with lower socioeconomic conditions [27]. It remains an open question whether LLM-based POI recommendation models can effectively generalize across diverse urban environments.
In addition, because many studies rely on the FSQ-NYC and FSQ-TKY, they are often constrained to the timespan it covers: check-in data from 2012 to 2013. However, POI data is inherently dynamic: venues may have closed, relocated, or changed in category over time. [61] underscores the importance of validating the temporal quality of POI datasets by recording whether and when a venue’s information has been updated to reflect real-world changes. This is particularly critical, as recommender systems should avoid suggesting POIs that no longer exist or have undergone substantial changes (e.g., a former bookstore converted into a coworking space) and behave dynamically over longitudinal periods [53]. Moreover, behavioral patterns captured over a decade ago may no longer align with modern user preferences and routines. For example, the opening of a new train station may significantly shift commuting patterns and the popularity of surrounding POIs.
@@
More recently, researchers have begun leveraging the broader Global-scale Check-in Dataset (GSCD) [55, 56], which spans 415 cities across 77 countries. Despite its wider geographic coverage, GSCD is temporally limited to the same 2012-2013 period as FSQ-NYC and FSQ-TKY, and thus suffers from similar issues of temporal quality. More critically, [29] demonstrated that GSCD suffers from significant data quality issues, with over 14 million check-ins (about 44%) of the dataset flagged as erroneous due to anomalous user behavior. These include (1) repeated check-ins at the same venue, (2) check-ins occurring within implausibly short time intervals (less than one minute), and (3) transitions between venues that would require travel speeds exceeding Mach 1, which are physically unreasonable.
To address these limitations, [29] introduced the Semantic Trails Dataset (STD), which applies systematic filtering procedures to enhance data quality. STD comprises two subsets: a cleaned version of GSCD covering 2012-2013 (STD 2013), and a newer collection of check-ins from 2017-2018 (STD 2018), sourced from Foursquare Swarm. STD 2018 also spans a wider range of cities, making it valuable for capturing globally distributed user behavior, in contrast to GSCD’s focus on densely populated urban centers. Both subsets follow the same rigorous filtering criteria, resulting in a higher-quality check-in dataset for POI recommendation tasks. Given these improvements, we adopted STD as the source for our check-in dataset.
@@
Another persistent challenge in POI recommendation research is the lack of reproducibility in dataset preprocessing. While some recent studies utilize datasets like GSCD to cover a wide range of cities, they often omit important details needed for replicating their data filtering processes. For example, AgentMove [7] conducted city-level filtering based on a minimum distance between each trajectory and the city center, but they did not specify how the city center was defined or what distance threshold was used. Generally, city filtering can be done in two ways: (1) by selecting points within a fixed radius from a central coordinate, or (2) by applying precise administrative boundaries, such as those from OpenStreetMap. The radius-based approach is less reliable, as it assumes a circular city shape and requires a subjective definition of the city’s center. In contrast, the boundary-based method is more robust and reproducible, and is the approach we adopted in this work. Without standardized, transparent filtering criteria and publicly available datasets, it remains difficult to ensure fair comparisons across POI recommendation methods.
@@ 2.2 Understanding Urban Features and POI Recommendation
POI recommendation studies that evaluate models across multiple city-level datasets often include analyses to assess how well their methods generalize across different urban contexts. It is well understood that POI recommendation accuracy metrics (e.g., Acc@k, NDCG@k) can vary substantially between cities and can be interpreted as a proxy for how easy or difficult a city is to model. The assumption is that higher performance reflects more predictable or structured mobility patterns. This viewpoint is consistent with prior work highlighting the role of cultural and urban-specific factors in shaping mobility behaviors [55, 34].
Several studies have proposed hypotheses connecting specific urban features to modeling difficulty. For example, GETNext [60] hypothesized that cities with fewer check-ins and higher spatial sparsity of POIs are harder to model. STHGCN [54] suggested that a larger number of user trajectories improves predictive accuracy by providing richer collaborative signals, whose architecture is designed to leverage. LLM4POI [22] proposed that cities with a greater variety of POI categories are easier to model due to LLMs’ contextual reasoning capabilities, whereas cities covering a broader geographic area tend to be more difficult to model. In the zero-shot POI recommendation setting, AgentMove  [7] reported two key findings: (1) geospatial biases inherent in LLMs can hinder prediction quality across cities, and (2) LLMs are influenced by city-specific mobility patterns.
@@ 3 Dataset
Table 2: Summary statistics of the 12 Massive-STEPS subsets, including the number of users, trajectories, POI locations, total check-ins, and train, validation, and test sample counts. For comparison, we also include statistics from existing Foursquare-based [57] and Gowalla-based [4] datasets. Due to variations in dataset preprocessing across studies, we report the version used in [54].
@@
NYC and Tokyo Check-in Dataset [57]
@@ 3.1 Creation Process
Massive-STEPS is derived from the two subsets of STD [29], incorporating check-ins from both the 2013 and 2018 subsets. We utilize two additional components from STD: (1) the cities metadata file, which provides the latitude and longitude of administrative regions (e.g., towns, suburbs) along with their corresponding country codes obtained from GeoNames; and (2) the POI category mapping, which links each Foursquare Category ID to its descriptive name (e.g., "Restaurant"). Based on this metadata, each POI is thus associated with several attributes: Foursquare Place ID, Foursquare Category ID, category name, latitude/longitude of the administrative region, the administrative region name, and the country code. For anonymization purposes and model training compatibility, we apply ordinal encoding to the Place IDs and Category IDs, assigning each a unique integer index.
@@ 3.1.1 Preprocessing
We split trajectories into training, validation, and test sets in a ratio of 7:1:2, following [7]. We ensure that all users in the test set appear at least once in the training or validation set, a common practice in prior studies [60, 54].
@@ 3.2 Description and Addressing Literature Gaps
Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) dataset is a city-level POI check-in dataset derived from Semantic Trails [29], comprising user check-in trajectories from 12 cities: Beijing, Istanbul, Jakarta, Kuwait City, Melbourne, Moscow, New York, Petaling Jaya, São Paulo, Shanghai, Sydney, and Tokyo. It features anonymized POI check-ins enriched with geographical metadata to support spatiotemporal and sequential modeling tasks. City-level statistics, along with comparisons to existing datasets, are presented in Table 2. Fig. 2 shows a world map highlighting the locations of all cities included in the Massive-STEPS dataset. Table 6 shows the available fields in the dataset and provides an example for each field.
Massive-STEPS offers a more comprehensive and diverse representation of urban mobility compared to typical POI check-in datasets. As shown in Table 2, datasets like FSQ-NYC and FSQ-TKY [57] contain fewer than 10,000 candidate POI locations. In contrast, cities in Massive-STEPS cover significantly more POIs: Massive-STEPS New York has over 49,000 POIs, while Massive-STEPS Jakarta exceeds 76,000. Massive-STEPS Istanbul, one of the largest subsets, features a large user base of 23,700, offering a broad range of user behaviors. Although some Massive-STEPS subsets are smaller than their FSQ counterparts (e.g., Tokyo), we attribute this to the strict filtering procedures applied by STD to remove erroneous entries, as explained in Section 2.1. This scale introduces additional computational challenges. Models that rely on dense POI-to-POI adjacency matrices, for instance, may require more efficient implementations to avoid excessive memory consumption.